OpenStudy (anonymous):

d-6 > -4 solve answers 1.d > -2 2.d > -10 3.d > 2 4.d >10

OpenStudy (anonymous):

3

OpenStudy (anonymous):

2-6=-4

OpenStudy (anonymous):

I have another

OpenStudy (anonymous):

12 + 4m =24

OpenStudy (anonymous):

answers 1. m =9 2.m=8 3.m=3 4.m=-6

OpenStudy (anonymous):

Get m by its self, subtract 12 from both side 4m = 24 - 12 4m = 12 Divide 4 from both sides m = 3

OpenStudy (anonymous):

thanks that makes sences

OpenStudy (anonymous):

another

OpenStudy (anonymous):

r/2 - 6 =14

OpenStudy (anonymous):

answers 1.r=40 2.r=34 3.r=20 4.r=16

OpenStudy (anonymous):

r=16

OpenStudy (anonymous):

r=40